OpenStudy (anonymous):

Evaluate (x + y)^0 for x = 4 and y = 3. A. 0 B. 1 C. 2 D. 7

Nnesha (nnesha):

x =4 so replace x by 4 and y by 3 and remember exponent rules!!

Nnesha (nnesha):

what would be your answer if there is 0 exponent ?

OpenStudy (anonymous):

(4+3) 7

OpenStudy (anonymous):

no 0

OpenStudy (acxbox22):

u dont really need to plug anything in since the expression is to the power of 0

Nnesha (nnesha):

again my question is \(\color{blue}{\text{Originally Posted by}}\) @Nnesha what would be your answer if there is 0 exponent ? \(\color{blue}{\text{End of Quote}}\)

OpenStudy (alexandervonhumboldt2):

a^0=1

OpenStudy (alexandervonhumboldt2):

let me prove:

OpenStudy (anonymous):

the answer is 7

OpenStudy (alexandervonhumboldt2):

a^0=a^1/a^1 a^1/a^1=a/a=1

OpenStudy (alexandervonhumboldt2):

the ansewer is not seven

OpenStudy (acxbox22):



Nnesha (nnesha):

nope @snickers23232 do you know exponent rules ?

OpenStudy (alexandervonhumboldt2):

anything in 0 exponent is 1

OpenStudy (alexandervonhumboldt2):

exept 0

OpenStudy (anonymous):

no =/

OpenStudy (anonymous):

so the answer is 1?

OpenStudy (alexandervonhumboldt2):

yes

Nnesha (nnesha):

whenever variable or number any number or variable raise to the 0 exponent answer always going to be 0